OpenStudy (anonymous):

Write the equation of the parabola with vertex (9, –2) and point (12, 16) in vertex form.

OpenStudy (anonymous):

general equation is y-h=a(x-b)^2 where b and h are coordinates of vortex (b,h) here b=9 , h=-2 so y-(-2)=a(x-9)^2 y+2=a(x-9)^2 y=a(x-9)^2-2 now put x=12 and y=16 and solve the equation to get value of a .

OpenStudy (anonymous):

A. y = (x + 9)2 + 2 B. y = 2(x – 12)2 – 16 C. y = 2(x – 9)2 – 2 D. y = 3(x + 12)2 + 16 I think it would be C

OpenStudy (anonymous):

smart guess !!!! yes it is .

OpenStudy (anonymous):

yes!